OpenStudy (anonymous):

when will an object with a slope of 3m/s be traveling at 34 m/s

OpenStudy (anonymous):

your question is not clear..

OpenStudy (anonymous):

ok like there is a graph that has a slope of 3m/s and the question is asking me when the object will be traveling at 34 m/s, like at 2 seconds had a velocity of 10 m/s, and at 5 sec had a velocity of 15 m/sec and so on

OpenStudy (shane_b):

I take it this is a velocity vs time graph so the slope at any given point equals the instantaneous acceleration (since dV/dt=a). \[\Delta V = at\]Rearrange that to get:\[\frac{\Delta V}{a}=t\]Now plug in values:\[\frac {34m/s}{3m/s^2}=11.333s\]

OpenStudy (ash2326):

@rcb95 you could see that from the graph, Can you post it?

OpenStudy (ash2326):

@Shane_B Velocity time graph's slope is m^2/s this is displacement time graph, that's why the slope is m/s

OpenStudy (shane_b):

I wish they'd just post the graph then :) From the wording I took a leap and decided it must be a V vs T graph

OpenStudy (anonymous):

Thanks so much @Shane_B

OpenStudy (anonymous):

rfl

OpenStudy (shane_b):

@rcb: It is a displacement vs time graph or a velocity vs time graph?

OpenStudy (shane_b):

@ash2326: I guess we'll never know :)